OpenStudy (babyslapmafro):

(Calc3, Lagrange multipliers) Please help me solve the following problem. Find the point on the line x+2y+z=1 that is closest to the origin.

OpenStudy (babyslapmafro):

Do I minimize x^2+y^2

OpenStudy (amistre64):

thats not a line, its a plane

OpenStudy (amistre64):

no, minimize a sphere at the origin

OpenStudy (babyslapmafro):

Oh right, typo

OpenStudy (amistre64):

f(x,y,z) = x^2 + y^2 + z^2

OpenStudy (babyslapmafro):

Ok and if it were a line then I would minimize x^2+y^2

OpenStudy (amistre64):

yes, distance from the origin to a line in the xy plane would represent a circle of x^2 + y^2

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (amistre64):

so double chk your results, define a line from the origin using the normal vector of the plane: x = t y = 2t z = t plug those into the plane equation to solve for t

OpenStudy (amistre64):

x+2y+z=1 t+2(2t)+t=1 6t=1, t = 1/6 the point should be (1/6, 1/3, 1/6)